Inventory records for Dunbar Incorporated revealed the following: Date Transaction Number of Units Unit Cost Apr. 1 Beginning inventory 470 $ 2.11 Apr. 20 Purchase 400 2.67 Dunbar sold 640 units of inventory during the month. Ending inventory assuming LIFO would be: (Do not round your intermediate calculations. Round your answer to the nearest dollar amount.) Multiple Choice $614. $485. $992. $1,068.
Answer #1

Ending inventory = (470+400) - 640 = 230 units

= (230 * 2.11)

= 485

Option B is the answer